A call option with a current value of $11.70. A put option with a current value of $7.70. Both options written on the same stock and both with 1 year until expiration. The current price of the stock is $48.00 and the prevailing risk-free rate is 9.00%. What must be the striking price of either option? *** In your calculations, use simple discounting instead of continuous discounting. Also, do not enter the dollar sign and use two decimals (round off to 2 decimals).
Answer #1

As per Put Call Parity Equation,

C + Xe-rt = P + S

11.70 + Xe-(0.09*1) = 7.70 + 48

X(0.9139) = 44

X = $48.15

So,

Exercise Price = $48.15